# Redistribution Through

### v2

@@ 3.3 Order-maximal segmentations and optimality
In segment $p^{\prime}$, including a mass $\delta$ of types $\tilde{\theta}$ increases the demand at any price $q\in\{p^{\prime},\dots,\tilde{\theta}\}$ by an amount $\delta$ while keeping the demand at any other price unaffected. Therefore, this transfer leads the seller to deviate in segment $p^{\prime}$ to some price $q$ with $p^{\prime}<q\leq\tilde{\theta}$ if and only if
$\delta>\frac{1}{q-p^{\prime}}\Biggl{(}p^{\prime}\sum_{\theta\geq p^{\prime}}\sigma(\theta,p^{\prime})-q\sum_{\theta\geq q}\sigma(\theta,p^{\prime})\Biggr{)}.$ (3)
In segment $p^{\prime\prime}$, excluding a mass $\delta$ of types $\tilde{\theta}$ decreases the demand at any price $q\in\{p^{\prime},\dots,\tilde{\theta}\}$ by an amount $\delta$. Therefore, the seller deems it profitable to deviate in segment $p^{\prime\prime}$ to some price $q$ with $q>\tilde{\theta}$ if and only if
$\delta>\frac{1}{p^{\prime\prime}}\Biggl{(}p^{\prime\prime}\sum_{\theta\geq p^{\prime\prime}}\sigma(\theta,p^{\prime\prime})-q\sum_{\theta\geq q}\sigma(\theta,p^{\prime\prime})\Biggr{)}.$ (4)
It is clear that no downward transfer of type $\tilde{\theta}$ from segment $p^{\prime\prime}$ to segment $p^{\prime}$ is feasible for $\sigma$ if, in one of the two segments $p^{\prime}$ and $p^{\prime\prime}$, at least one of the obedience constraints relative to the deviations highlighted above is binding. Indeed, this would imply that the term on the right-hand side of either inequality 3 or 4 is zero.
We now turn to the conditions under which a particular redistributive transfer is unfeasible for some segmentation. Let $\mu\in\Delta(\Theta)$ and $\sigma\in\Sigma_{\Omega}(\mu)$, and assume there exists $p^{\prime},p^{\prime\prime}\in\operatorname{supp}(\sigma_{P})$, $\theta^{\prime}\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p^{\prime\prime})\bigr{)}$ and $\theta^{\prime\prime}\in\operatorname{supp}\bigl{(}\sigma(\,\cdot\,|\,p^{\prime})\bigr{)}$ with $\theta^{\prime\prime}>\theta^{\prime}\geq p^{\prime\prime}>p^{\prime}$. Consider a redistributive buyer transfer such that a mass $\varepsilon$ of buyers of type $\theta^{\prime}$ are transferred from segment $p^{\prime\prime}$ to $p^{\prime}$, while the same mass $\varepsilon$ of buyers of type $\theta^{\prime\prime}$ are transferred from segment $p^{\prime}$ to $p^{\prime\prime}$, as illustrated in Figure 3(b). This transfer does not impact obedience constraints in segment $p^{\prime}$. Indeed, it decreases the demand at any price $\theta^{\prime}<q\leq\theta^{\prime\prime}$ and does not impact the demand at any other price. In segment $p^{\prime\prime}$, however, the transfer increases the demand at any price $q$ such that $\theta^{\prime}<q\leq\theta^{\prime\prime}$, while leaving the demand at any other price unchanged. Therefore, the redistributive transfer prompts the seller to deviate in segment $p^{\prime\prime}$ to charging a price $\theta^{\prime}<q\leq\theta^{\prime\prime}$ if and only if
$\varepsilon>\frac{1}{q}\Biggl{(}p^{\prime\prime}\sum_{\theta\geq p^{\prime\prime}}\sigma(\theta,p^{\prime\prime})-q\sum_{\theta\geq q}\sigma(\theta,p^{\prime\prime})\Biggr{)}.$ (5)
Hence, no redistributive transfer between types $\theta^{\prime}$ and $\theta^{\prime\prime}$ in segments $p^{\prime}$ and $p^{\prime\prime}$ is feasible for $\sigma$ if, in the upper segment $p^{\prime\prime}$, at least one of the obedience constraints relative to the deviations highlighted above is binding, as this would imply that the term on the right-hand side of inequality 5 is zero.
Proposition 2 establishes necessary and sufficient conditions a segmentation must satisfy for no downward or redistributive transfers to be feasible for that segmentation. To see why, note that condition (ii) guarantees that no buyer of type $\theta$ belonging to some segment $p$ can be transferred to some segment $p^{\prime}$ such that $p<p^{\prime}<\theta$ because the obedience constraint relative to the price $\theta$ is binding in segment $p^{\prime}$. Since redistributive transfers always involve an upward transfer of this sort, condition (ii) ensures that one cannot perform any such transfer. [^21] An upward transfer is defined as the opposite of a downward transfer (Definition 2). Condition (ii) also ensures that some downward transfers are not feasible. Buyers of type $\theta$ cannot be transferred from a segment $p^{\prime}$ to any lower segment $p$ which includes some buyers of type $\theta^{\prime}>\theta$ because the obedience constraint for the price $\theta^{\prime}$ must be binding in segment $p^{\prime}$. However, a downward transfer could in principle be possible if the lower segment $p$ did not include any buyer of type $\theta^{\prime}>\theta$. Condition (i) blocks all these remaining downward transfers. It ensures that even if no buyer of type $\theta^{\prime}>\theta$ is assigned to segment $p$, the obedience constraint relative to some price $\tilde{\theta}\leq\theta$ must be binding in segment $p$, which renders such downward transfer unfeasible.
3.4 Monotone segmentations and progressive pricing
@@
In this subsection, we show that redistributive segmentations exhibit a positive-assortative structure. In other words, redistributive segmentations assign low types to low-price segments and high types to high-price segments. We introduce two notions of positive-assortativity for segmentations.
@@
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is weakly monotone if for any $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$,
@@
Definition 5 (Strongly monotone segmentations).
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ is strongly monotone if for any $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$,
Figures 4(a) and 4(b) respectively illustrate weakly and strongly monotone segmentations.
$\theta$$p$
@@
A segmentation is weakly monotone if, for any pair of segments, the highest type assigned to the lower-price segment is not strictly greater than the highest type assigned to the higher-price segment. Whereas a segmentation is strongly monotone if, for any pair of segments, all buyers in the lower-price segment have (weakly) lower types than all buyers in the higher-price segment. [^22] Note that our definition of weak monotonicity implies that the segment supports are weak-set-ordered (Che et al., 2021). Strong monotonicity, on the other hand, implies an ordering of segment supports that is stronger than the strong set order (Topkis, 1998). Note also that, since the intersection of any pair of segment supports contains at most one element, strongly monotone segmentations correspond to a slightly weaker version of monotone partitional information structures (Ivanov, 2021; Mensch, 2021; Onuchic and Ray, 2023; Kolotilin et al., 2024) which partition the state space into convex sets.
Both concepts of monotonicity relate to some form of progressive pricing where high types are charged higher prices than low types. In weakly monotone segmentations, for every buyer of a given type, there exists a buyer of a higher type that is charged a (weakly) higher price. Nevertheless, this notion of progressivity is mild since some buyers of lower type might pay higher prices than some buyers of higher type. This outcome arises in regions where the support of different segments overlap, as buyer types belonging to such overlap are randomly assigned to different segments. On the other hand, strongly monotone segmentations are purely progressive. If a type is consuming at a given price, then every higher type must be assigned to a (weakly) higher segment. Indeed, strongly monotone segmentations assign buyers to segments deterministically, except perhaps for types located at the intersection of two consecutive segments.
A direct implication of condition (ii) in Proposition 2 is that redistributive segmentations are weakly monotone. Indeed, for any $\succsim_{\mathcal{R}}$-maximal segmentation, if a type $\theta$ is in a segment $p\in\operatorname{supp}(\sigma_{P})$, then it must belong to any other segment $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}\leq\theta$, which implies weak monotonicity. [^23] We state the result for strictly redistributive welfare functions to clarify the exposition. If $w\in\mathcal{R}$, then for any $\mu\in\Delta(\Theta)$, there exists a $\mathrm{P}^{\star}_{w,\mu}$-optimal and weakly monotone segmentation of $\mu$.
Theorem 2.A.
@@
Strongly monotone segmentations, which represent pure progressive pricing, imply that all possible redistributive transfers have been exhausted. This raises the question: under what conditions on redistributive welfare functions do we wish to exhaust all opportunities for implementing such transfers? Suppose we aim to perform a redistributive transfer on a segmentation that is not strongly monotone. While these transfers are always desirable from a welfare perspective, they may not be feasible because they could violate obedience constraints in the upper segment (cf. Figure 3(b)). Nevertheless, it is possible to preserve obedience in the upper segment by transferring a certain mass of the highest type in this segment to a higher-price segment. This series of transfers improves aggregate welfare only if the net welfare gain from the redistributive transfer exceeds the welfare loss due to this upward transfer. The following concept of strongly redistributive welfare functions precisely captures the necessary and sufficient condition for this series of transfers to always be beneficial.
@@
A welfare function is strongly redistributive if the net welfare gain from a redistributive transfer involving the two lowest types that could be included in the segment $\theta_{k}$—given by the expression on the left-hand side of condition 6—exceeds by a factor $\theta_{k+1}/(\theta_{k+1}-\theta_{k})$ the maximal welfare loss induced by the upward transfer of any higher type $\theta_{\ell}$—given by the second term on the right-hand side. Figure 5 illustrates the two transfers involved in 6.
$\theta$$p$$p$$\theta_{k}$$\theta_{k}$$\theta_{k+1}$$\theta_{\ell}$$\varepsilon$$\varepsilon$$\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\,\varepsilon$
Figure 5: The combination of a redistributive transfer (in orange) and an upward transfer of high-types in the upper segment (in blue).
@@
As the next result shows, strongly redistributive welfare functions are the only functions for which every optimal segmentation is strongly monotone. We explain condition 6 together with the theorem below.
Theorem 2.B.
@@
A welfare function $w$ is strongly redistributive if and only if for any $\mu\in\Delta(\Theta)$, every $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is strongly monotone.
The proof of Theorem 2.B can be found in Appendix D. In condition 6, $\theta_{k}$ is both the type that benefits from the redistributive transfer and the price charged at the upper segment involved in this transfer. Focusing on such transfers is sufficient because whenever we can implement a redistributive transfer, we can always ensure that it benefits the type equal to the price in the upper segment. Conversely, to exhaust all possible redistributive transfers, it is necessary to perform those that benefit the type corresponding to the price of the upper segment involved in the transfer.
The second term on the right-hand side of 6 corresponds to the maximum welfare loss the type $\theta_{\ell}$ can incur by being transferred upwards from segment $\theta_{k}$. Indeed, the segment $\theta_{\ell}$ is the highest one where the type $\theta_{\ell}$ can be transferred while ensuring efficiency. This welfare loss is scaled by the ratio $\theta_{k+1}/(\theta_{k+1}-\theta_{k})>1$. The reason is the following. If we implement a redistributive transfer of a mass $\varepsilon$ of types $\theta_{k}$ and $\theta_{j}>\theta_{k}$, and this transfer breaks Ob, then the seller will deviate to charging a price $\theta_{k}<p\leq\theta_{j}$ in the upper segment (cf. Figure 3(b)). To preserve obedience, an even larger mass of higher types $\theta_{\ell}\geq\theta_{j}$ must be transferred upward from the upper segment. Specifically, a mass $\varepsilon\times p/(p-\theta_{k})$ of types $\theta_{\ell}$ must be transferred. The ratio $p/(p-\theta_{k})$ decreases with $p$ and is thus maximal for $p=\theta_{k+1}$.
The term on the left-hand side of 6 corresponds to the net welfare gain from a redistributive transfer between two consecutive types. By supermodularity of $w$, the net welfare gain from a redistributive transfer between $\theta_{k}$ and $\theta_{k+1}$ is smaller than the net gain from a redistributive transfer between $\theta_{k}$ and any $\theta_{j}$, with $j>k+1$. As such, a redistributive transfer between these two consecutive types yields the minimal net welfare gain from a redistributive transfer benefiting type $\theta_{k}$.
To summarize, satisfying the inequality 6 implies that the minimal net welfare gains from performing a redistributive transfer benefiting type $\theta_{k}$ outweighs the maximal welfare loss from performing an upward transfer that involves the maximal mass of types $\theta_{\ell}$. Therefore, it is both necessary and sufficient to make any opportunity to implement a redistributive transfer desirable.
We illustrate strongly redistributive welfare functions with Examples 1 and 2. This shows incidentally that the set of these functions is non-empty.
@@
Let $\lambda:\Theta\to\mathbb{R}$ be a strictly decreasing function such that $\lambda(\theta)>0$ for all $\theta\in\Theta$. Define $w$ such that, for any $(\theta,p)\in\Omega$, $w(\theta,p)=\lambda(\theta)(\theta-p)$. By rearranging terms in the inequality 6, we can see that $w$ is strongly redistributive if and only if for any $\theta_{k},\theta_{\ell}\in\Theta$ such that $\theta_{k}<\theta_{\ell}$,
$\frac{\lambda(\theta_{k+1})}{\lambda(\theta_{\ell})}\left(\frac{\lambda(\theta_{k})}{\lambda(\theta_{k+1})}-1\right)\geq\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\frac{\theta_{\ell}-\theta_{k}}{\theta_{k}-\theta_{k-1}}.$ (7)
Inequality 7 imposes a condition on the rate of decrease of $\theta\mapsto\lambda(\theta)$. In this case, we can reinterpret Theorem 2.B as follows: every optimal segmentation is strongly monotone if and only if Pareto weights decrease at a sufficiently fast rate.
@@
Let $u:\mathbb{R}\to\mathbb{R}$ be strictly increasing and strictly concave. Define $w$ such that, for any $(\theta,p)\in\Omega$, $w(\theta,p)=u(\theta-p)$. In this case, one can show that condition 6 can never be satisfied. Conversely to Example 1, when the redistributive concerns stem from a concave transformation of buyers’ surplus, the welfare gain associated with redistributive transfers can never outweigh the welfare loss associated with burning the surplus of some high type. In this case, Theorem 2.B implies that there always exists a market $\mu\in\Delta(\Theta)$ such that no optimal segmentation is strongly monotone.
Theorem 2.B establishes that a welfare function is strongly redistributive if and only if any optimal segmentation is strongly monotone. Furthermore, as we show in the following lemma (which proof can be found in Appendix E), for any given aggregate market, there is only one strongly monotone segmentation of that market that is $\succsim_{\mathcal{R}}$-maximal. This particular segmentation therefore represents the uniquely optimal market segmentation for any strongly redistributive welfare function.
Lemma 2.
For any $\mu\in\Delta(\Theta)$, there exists a unique strongly monotone segmentation $\sigma\in\Sigma_{\Omega}(\mu)$ that is $\succsim_{\mathcal{R}}$-maximal.

### v5

@@ 3.2.1. Intuitive Explanation of Theorem˜1
This transfer affects the seller’s obedience constraints in both segments. In the lower segment $p^{\prime}$, adding these buyers increases demand at prices between $p^{\prime}$ and $\tilde{\theta}$, tightening constraints for these prices (shown in the orange box in the lower segment). Simultaneously, in the higher segment $p^{\prime\prime}$, removing buyers decreases demand at prices between $p^{\prime\prime}$ and $\tilde{\theta}$, tightening constraints for prices above $\tilde{\theta}$ (represented in the orange box in the upper segment). If at least one of the obedience constraints is already binding at those prices in either segment, even a small transfer will violate the seller’s obedience constraints, making the downward transfer infeasible.
The reasoning is similar for redistributive transfers. Consider swapping some mass of buyers between segments as shown in Figure˜3(b): buyers of type $\theta^{\prime}$ move from segment $p^{\prime\prime}$ to $p^{\prime}$, while buyers of type $\theta^{\prime\prime}$ move from $p^{\prime}$ to $p^{\prime\prime}$. This transfer does not tighten any of the obedience constraints in segment $p^{\prime}$. Indeed, it decreases the demand at all prices between $\theta^{\prime}$ and $\theta^{\prime\prime}$, and does not affect the demand at any other price. However, in the higher segment $p^{\prime\prime}$, it increases demand at all these intermediate prices (highlighted by the orange box in Figure˜3(b)), potentially violating obedience constraints in that segment.
Saturated segmentations are precisely those for which no downward or redistributive transfers remain feasible. Condition (ii) ensures that no buyer of type $\theta$ belonging to segment $p$ can be moved upwards, as the obedience constraint for price $\theta$ binds in every upper segment they could be moved to. [^22] Since we are restricting attention to efficient segmentations, this corresponds to segments $p^{\prime}$ such that $p<p^{\prime}\leq\theta$. This blocks all redistributive transfers, which necessarily involve moving some buyers upward. Condition (ii) also blocks some downward transfers: one cannot transfer some type $\theta$ downwards to a segment that includes any higher type $\theta^{\prime}>\theta$, since this would lead to a violation of obedience in the origin segment. Transferring $\theta$ downwards to a segment without any higher types, however, could in principle be possible. Condition (i) blocks those remaining downward transfers by ensuring that even if such transfer does not lead to a violation of obedience in the origin segment, it does in the destination one.
3.2.2. The Proof of Theorem˜1
The proof of Theorem˜1 is based on a key result of independent interest. We articulate a sense in which one segmentation is more redistributive than another by introducing a partial order on the space of efficient segmentations (Definition˜2). Under this order, a segmentation is more redistributive than another if the former generates higher aggregate welfare than the latter for every redistributive welfare function. Proposition˜1 establishes that this occurs precisely when the more redistributive segmentation can be obtained through a sequence of downward and redistributive buyer transfers applied to the less redistributive one. [^23] This provides an answer to a practical question: Say one observes some market segmentation. How can it be modified to make it more redistributive? We then demonstrate that optimal redistributive segmentations are exactly those for which no segmentation that is more redistributive exists (Lemma˜2) and that this condition is equivalent to being saturated (Lemma˜3).
Proof of Theorem˜1.
We first introduce a redistributive order on the space of efficient segmentations: [^24] See Section 3.5.2 for the detailed connection between our redistributive order and existing stochastic orders, in particular, the supermodular one (shaked2007; Muller2013; Meyer2015).
Definition 2 (Redistributive order).
Let $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$. We say that $\sigma$ is more redistributive than $\sigma^{\prime}$, denoted $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$, if
$\forall w\in\mathcal{R},\quad\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma(\theta,p)\geq\sum_{(\theta,p)\in\Theta\times P}w(\theta,p)\,\sigma^{\prime}(\theta,p).$
[…]
We call a buyer mass transfer (in short, a transfer) any $t\colon\Omega\to\mathbb{R}$ that satisfies $\sum_{p\leq\theta}t(\theta,p)=0$ for any $\theta\in\Theta$. Intuitively, transfers perturb a segmentation $\sigma$ by moving buyers between segments, with $t(\theta,p)$ denoting the mass of buyers of type $\theta$ that is added to or removed from a segment with price $p$. We now define two particular classes of transfers.
Definition 3 (Downward transfers).
A transfer $t$ is downward if there exist $\delta\geq 0$ and $\tilde{\theta},p^{\prime},p^{\prime\prime}\in\Theta$ such that $p^{\prime}<p^{\prime\prime}$, and
Definition 4 (Redistributive transfers).
A transfer $t$ is redistributive if there exists $\varepsilon\geq 0$ and $(\theta^{\prime},p^{\prime\prime}),(\theta^{\prime\prime},p^{\prime})\in\Omega$ such that $\theta^{\prime}<\theta^{\prime\prime}$, $p^{\prime}<p^{\prime\prime}$, and
Proposition˜1 shows that performing a sequence of such transfers is not only sufficient to make a segmentation more redistributive (according to Definition˜2) but also necessary. Let $\mathcal{T}$ be the set of all downward and redistributive transfers.
Proposition 1.
Let $\mu\in\Delta(\Theta)$ and $\sigma,\sigma^{\prime}\in\Sigma^{\mathrm{eff}}(\mu)$. The two following claims are equivalent:
  1. (i)
    $\sigma\succsim_{\mathcal{R}}\sigma^{\prime}$;
  2. (ii)
    There exists $n\in\mathbb{N}$, $n>0$, and $(t_{i})_{i\in\{1,\dots,n\}}\in\mathcal{T}^{n}$ such that
    $\sigma=\sigma^{\prime}+\sum_{i=1}^{n}t_{i}.$
The proof of Proposition˜1, which can be found in Appendix˜A, relies on duality arguments akin to those employed by Muller2013 and Meyer2015. Specifically, we show that the space of redistributive welfare functions (up to an additive constant) is dual to the cone of downward and redistributive transfers.
[…]
Lemma 2.
For any $\mu\in\Delta(\Theta)$, the following two claims are true:
  1. (i)
[…]
  2. (ii)
[…]
Lemma 3.
[…]
$3.$  Monotonicity and Progressive Pricing
@@ 3.2.2. The Proof of Theorem˜1
In this subsection, we show an important economic implication of Theorem˜1. Optimal redistributive segmentations exhibit a positive assortative structure. They assign low-type buyers to low-price segments and high-type buyers to high-price segments. Furthermore, this positive-assortative structure is strenghtened as redistributive concerns grow.
3.3.1. Weak Monotonicity
We start by introducing a weak notion of positive assortativity for segmentations:
@@ 3.3.1. Weak Monotonicity
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is weakly monotone if for any $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$,
@@
[…]
@@
A segmentation is weakly monotone if, for any pair of segments, the highest type assigned to the lower-price segment is not strictly greater than the highest type assigned to the higher-price segment. Figure˜4(a) illustrates weakly monotone segmentations.
A direct implication of Theorem˜1 is that all optimal redistributive segmentations are weakly monotone. Indeed, by condition (b) in Definition˜1, for any saturated segmentation, if a type $\theta$ is in a segment $p\in\operatorname{supp}(\sigma_{P})$, then it must belong to any other segment $p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}\leq\theta$, which implies weak monotonicity. [^26] We state the result for strictly redistributive welfare functions to clarify the exposition. If $w\in\mathcal{R}$, then for any $\mu\in\Delta(\Theta)$, there exists a $\mathrm{P}^{\star}_{w,\mu}$-optimal and weakly monotone segmentation of $\mu$.
Corollary 1.
@@
Corollary˜1 establishes that redistributive segmentations induce a form of progressive pricing, whereby high-types are charged higher prices than low-types. In weakly monotone segmentations, for every buyer of a given type, there exists a buyer of a higher type that is charged a (weakly) higher price. This notion of progressivity is mild, since some buyers of a lower type might still pay higher prices than some buyers of higher type. This outcome arises in regions where the support of different segments overlap, i.e., when buyers are randomly assigned to segments.
[…]
(a) A saturated weakly monotone segmentation.
[…]
(b) A compensated redistributive transfer.
Figure 5: The left panel illustrates a saturated segmentation where some buyers $\theta_{k+1}$ consume at a lower price than some buyers $\theta_{k}$. Points in orange represent $(\theta,p)$ for which Ob binds.
3.3.2. Strengthening Monotonicity
Compensated redistributive transfers
Regions of random assignment are exactly where unresolved redistributive tensions remain in saturated segmentations. Consider the segmentation shown in Figure˜5(a): some buyers of type $\theta_{k}$ are placed in the higher segment to “make room” for some buyers of type $\theta_{k+1}$ in the lower segment. These $\theta_{k}$ buyers would be better off in the lower segment, but in a saturated segmentation this cannot be achieved by simple downward or redistributive transfers without violating Ob. In particular, any redistributive transfer between types $\theta_{k}$ and $\theta_{k+1}$ would induce the seller to deviate to price $\theta_{k+1}$ in the upper segment.
Yet such a reallocation of types $\theta_{k}$ to the lower segment can be feasibly implemented through an appropriate sequence of transfers. Consider the construction illustrated in Figure˜5(b): first, a redistributive transfer swaps a mass $\varepsilon$ of types $\theta_{k}$ and $\theta_{k+1}$; then an upward transfer moves a mass $\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\varepsilon$ of higher-type buyers $\theta_{\ell}$ to a segment where they are charged exactly their willingness to pay. Removing this mass of type $\theta_{\ell}$ exactly offsets the effect of adding an $\varepsilon$ mass of types $\theta_{k+1}$ to segment $\theta_{k}$, preserving obedience. Moreover, because the transferred $\theta_{\ell}$ buyers are the highest types in segment $\theta_{k}$ and are moved to $(\theta_{\ell},\theta_{\ell})$, Ob remains satisfied. Thus, whenever a segmentation contains multiple types at the overlap between segments, one can construct such a feasible combination of redistributive and upward transfers. We refer to this construction as compensated redistributive transfers.
Strongly redistributive welfare functions
Compensated redistributive transfers may be desirable under welfare functions that reflect strong redistributive concerns. That is, welfare functions under which the welfare gain from lowering the price paid by a buyer $\theta_{k}$ offsets not only the welfare loss from an equivalent price increase for a buyer $\theta_{k+1}$ but also potential welfare losses borne by other higher types. Specifically, the compensated redistributive transfer illustrated in Figure˜5 is beneficial whenever the net welfare gain from the redistributive transfer between types $\theta_{k}$ and $\theta_{k+1}$ exceeds by a factor of $\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}$ the welfare loss from raising the price paid by a buyer of type $\theta_{\ell}$ from $\theta_{k}$ to $\theta_{\ell}$.
We now define a class of welfare functions for which every compensated redistributive transfer is desirable:
@@ 3.3.2. Strengthening Monotonicity
Strong monotonicity
Compensated redistributive transfers render any segmentation more positive-assortative by removing cases where a higher-type buyer pays less than a lower-type buyer. [^27] Note that performing a compensated redistributive transfer on a saturated segmentation also increases the seller’s profit, since the upward transfer increases the price being charged to $\theta_{\ell}$ while the redistributive transfer is profit-neutral. This aspect of compensated redistributive transfers is further examined in Section 3.4. As redistributive concerns become stronger, more compensated redistributive transfers should be performed. This process reaches a limit once no further compensated redistributive transfers can be implemented, yielding what we call strongly monotone segmentations:
Definition 7 (Strongly monotone segmentations).
For any $\mu\in\Delta(\Theta)$, a segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ is strongly monotone if for any $p,p^{\prime}\in\operatorname{supp}(\sigma_{P})$ such that $p<p^{\prime}$,
A segmentation is strongly monotone if buyers are assigned to segments deterministically, except possibly for types lying at the boundary between two consecutive segments. [^28] Our definition of weak monotonicity implies that segment supports are weak-set-ordered (Che2021). Strong monotonicity, by contrast, induces an ordering of segment supports stronger than the strong set order (topkis). Since the intersection of any two segment supports contains at most one element, strongly monotone segmentations correspond to a slightly weaker notion than monotone partitional information structures, which partition the state space into convex sets (Ivanov2021; Mensch2021; Kolotilin2022; Onuchic2023; Kolotilin2024). Strongly monotone segmentations, illustrated in Figure˜4(b), thus represent a pure form of progressive pricing: for any type, every higher type is charged a (weakly) higher price.
Optimality of strong monotonicity
@@
Our next result shows that strongly redistributive welfare functions are the only ones for which every optimal segmentation is strongly monotone.
Proposition 2.
@@
Any $\mathrm{P}^{\star}_{w,\mu}$-optimal segmentation is strongly monotone for every $\mu\in\Delta(\Theta)$ if and only if $w\in\overline{\mathcal{R}}$ is strongly redistributive.
The proof of Proposition˜2 appears in Appendix˜D. We further show in the following lemma (the proof of which can be found in Appendix˜E) that, for any given aggregate market, there is exactly one strongly monotone segmentation of that market that is saturated. This particular segmentation therefore represents the uniquely optimal market segmentation for any strongly redistributive welfare function.
Lemma 4.
For any $\mu\in\Delta(\Theta)$, there exists a unique strongly monotone segmentation $\sigma\in\Sigma^{\mathrm{eff}}(\mu)$ that is saturated.
We now come back to Examples˜1 and 2. We show that the weighted consumer surplus welfare function is strongly redistributive if and only if Pareto weights decrease sufficiently fast in willingess to pay. [^29] This shows incidentally that the set of strongly redistributive welfare functions is non-empty. We also show that concave increasing transformations of consumer surplus are never strongly redistributive.
@@
Let $\lambda\colon\Theta\to\mathbb{R}$ be a strictly decreasing function such that $\lambda(\theta)>0$ for all $\theta\in\Theta$. Define $w$ such that, for any $(\theta,p)\in\Omega$, $w(\theta,p)=\lambda(\theta)(\theta-p)$. By rearranging terms in the inequality 3, we can see that $w$ is strongly redistributive if and only if for any $\theta_{k},\theta_{\ell}\in\Theta$ such that $2\leq k\leq K-1$ and $k<\ell\leq K$,
$\frac{\lambda(\theta_{k})-\lambda(\theta_{k+1})}{\lambda(\theta_{\ell})}\geq\frac{\theta_{k+1}}{\theta_{k+1}-\theta_{k}}\frac{\theta_{\ell}-\theta_{k}}{\theta_{k}-\theta_{k-1}}.$ (4)
Inequality 4 imposes a condition on the rate of decrease of $\theta\mapsto\lambda(\theta)$. In this case, we can reinterpret Proposition˜2 as follows: every optimal segmentation is strongly monotone if and only if Pareto weights decrease at a sufficiently fast rate. For instance, weights of the form $\lambda(\theta)=\theta^{-\gamma}$ or $\lambda(\theta)=\exp(-\gamma\theta)$ are strongly redistributive if $\gamma$ exceeds some threshold.
@@
Let $u:\mathbb{R}\to\mathbb{R}$ be strictly increasing and strictly concave. Define $w$ such that, for any $(\theta,p)\in\Omega$, $w(\theta,p)=u(\theta-p)$. In this case, one can show that condition 3 can never be satisfied. In contrast to Example˜1, when the redistributive concerns stem from a concave transformation of buyers’ surplus, the welfare gain associated with redistributive transfers can never outweigh the welfare loss associated with burning the surplus of some high type. In this case, Proposition˜2 implies that there always exists a market $\mu\in\Delta(\Theta)$ such that no optimal segmentation is strongly monotone.